OpenStudy (anonymous):

8-3x=7x+8..what and how do i solve this?

OpenStudy (anonymous):

8-3x=7x+8 Since 7x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 7x from both sides. -3x+8-7x=8 Since -3x and -7x are like terms, subtract 7x from -3x to get -10x. -10x+8=8 Since 8 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 8 from both sides. -10x=-8+8 Add 8 to -8 to get 0. -10x=0 Divide each term in the equation by -10. -(10x)/(-10)=(0)/(-10) Simplify the left-hand side of the equation by canceling the common factors. x=(0)/(-10) Any expression with zero in the numerator is zero. x=0

OpenStudy (anonymous):

Thank you!